Sample ETOPIM article for Physica B: Condensed Matter

A. R. McGurn^{a,1}, P. Bhattacharya^{b}, J. Sabarinathan^{b}, W-D. Zhou^{b}, and P-C. Yu^{b}

a Department of Physics Western Michigan University Kalamazoo, MI 49008 USA

b Department of Electrical Engineering and Computer Science University of Michigan Ann Arbor, MI 48109-2122 USA

Abstract

Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.

Key words: Xxx. Xxx. Xxx. Xxx.

Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.

Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx

1 Correspondence: Prof. Arthur R. McGurn, Department of Physics, Western Michigan University, Kalamazoo, MI 49008 USA FAX: XXX-XXX-XXX

email mcgurn@wmich.edu

2 This work supported by Army Research Office Grant DAAD XXX-XXX-XXX and Office of Naval Research Grant NNN-NNN-NNN.

Preprint submitted to Elsevier Preprint 21 June 2002 Fig. 1. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.

xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.

Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.

Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.

Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.

In Fig. 1, xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.

Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.

Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx

Fig. 2. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.

xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.

Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.

Fig. 2 shows Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.

References

[1] Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.

[2] Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.

[3] Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx. Xxx xxx xxx xxx.